OpenStudy (anonymous):

What is the slope of the line that passes through the pair of points? (1, 7), (10, 1)

OpenStudy (ghazi):

slope= (1-7)/(10-1)

OpenStudy (ghazi):

that is (y2-y1)/(x2-x1)

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (babyslapmafro):

(y2-y1)/(x2-x1)=slope

OpenStudy (anonymous):

so is the answer 2/3 ??

OpenStudy (ghazi):

no

OpenStudy (anonymous):

close \[\frac{1-7}{10-1}=\frac{-6}{9}=-\frac{2}{3}\]

OpenStudy (anonymous):

-2/3

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

yes yuh i do this !!! hahah